1.   Each main bearing cap in an engine contains 4 bolts. The bolts are selected at random without replacement from a parts bin that contains 30 bolts from one supplier and 70 bolts from another.

a.    

b.   What is the probability that a main bearing cap contains all bolts from the same supplier?

 

 

c.   What is the probability that exactly 3 bolts are from the same supplier?


1
Expert's answer
2022-02-25T04:16:56-0500

Let "A" denote the event "bolt from the first supplier". Let "B" denote the event "bolt from the second supplier".

There are "\dbinom{30+70}{4}=3921225" possible outcomes.

a. The probability that the number of bolts from each supplier is the same is


"P(2\ same \ \&\ 2\ other)=P(2A, 2B)"

"=\dfrac{\dbinom{30}{2}\dbinom{70}{2}}{\dbinom{100}{4}}=\dfrac{435(2415)}{3921225}=0.267907"

b.  The probability that a main bearing cap contains all bolts from the same supplier is


"P(4\ same)=P(4A, 0B)+P(0A, 4B)"

"=\dfrac{\dbinom{30}{4}\dbinom{70}{0}}{\dbinom{100}{4}}+\dfrac{\dbinom{30}{0}\dbinom{70}{4}}{\dbinom{100}{4}}"

"=\\dfrac{\t27405(1)+1(\t916895)}{3921225}=0.240818"

c. the probability that exactly 3 bolts are from the same supplier


"P(3\ same \ \&\ 1\ other)=P(3A, 1B)+P(1A, 3B)"

"=\dfrac{\dbinom{30}{3}\dbinom{70}{1}}{\dbinom{100}{4}}+\dfrac{\dbinom{30}{1}\dbinom{70}{3}}{\dbinom{100}{4}}"

"=\dfrac{4060(70)+30(54740)}{3921225}=0.491275"